1. d) A disc, starting from rest, undergoes a constant angular acceleration for a time ti and reaches an angular velocity of w1. It then undergoes a constant negative acceleration and comes to rest in a time tz. Find an expression in terms of the quantities ti, tz, and wi for the total number of revolutions of the disc in the
Answer #1

Step 1: Find revolutions made by disk when speeding up

Using 1st rotational kinematic equation:

w1 = w0 + \alpha_1 *t1

Given that disk starts from rest, So w0 = 0 rad/sec

w1 = 0 + \alpha_1 *t1

\alpha_1 = w1/t1 = angular acceleration of disk when speeding up

Now Using 2nd rotational kinematic equation:

\theta_1 = w0*t1 + (1/2)*\alpha_1*t1^2

\theta_1 = 0*t1 + (1/2)*(w1/t1)*t1^2

\theta_1 = (w1*t1)/2

Step 2: Now find number of revolutions when it undergoes negative constant angular acceleration

Using 1st rotational kinematic equation:

w2 = w1 - \alpha_2 *t2

Given that disk finally stops, So w2 = 0 rad/sec

0 = w1 - \alpha_2 *t2

\alpha_2 = w1/t2 = angular de-acceleration of disk when speeding down

Now Using 2nd rotational kinematic equation:

\theta_2 = w1*t2 - (1/2)*\alpha_2*t2^2

\theta_2 = w1*t2 - (1/2)*(w1/t2)*t2^2

\theta_2 = (w1*t2)/2

So total angular displacement of disks will be:

\theta = \theta_1 + \theta_2

\theta = (w1*t1)/2 + (w1*t2)/2

\theta = w1*(t1 + t2)/2

Let me know if you've any query.